# On the Brun–Titchmarsh theorem. I

### v3

@@ 7.3. Estimates for exponential sums in $\mathcal{R}^{*}_{2}$
×$\ \ \ \ \times\sum_{1\leqslant|h|\leqslant H}\widetilde{F}\Big{(}\frac{h}{m_{2}n_{2}\delta};m_{1}n_{1},m_{2}n_{2},r\Big{)}\mathrm{e}\Big{(}\frac{(q/\delta)hr\overline{m_{1}n_{1}}}{m_{2}n_{2}}\Big{)}.$
@@
We would like explore the cancellations among the exponentials $\mathrm{e}(\frac{(q/\delta)hr\overline{m_{1}n_{1}}}{m_{2}n_{2}})$ using Lemma 7.2. It is now necessary to separate variables. By Fourier inversion,
$\ \ \ \ \widetilde{F}\Big{(}\frac{h}{m_{2}n_{2}\delta};m_{1}n_{1},m_{2}n_{2},r\Big{)}$
@@
Note that for all $\nu_{1},\nu_{2}\in\mathbf{N},$
We need further to separate variables in $\widehat{F}(*,*)$ by Mellin inversion. To this end, we truncate the integral over $y_{1}$ over five segments
@@
which is also negligible. It thus suffices to treat the integral over $y_{1}\in I_{2}\cup I_{4}$. In a similar manner, we may also reduce the integral over $y_{2}$ to $[-V_{2},-U]\cup[U,V_{2}]$, where
@@
By symmetry, we only transform the integrals over $y_{1}\in[U,V_{1}]$ and $y_{2}\in[U,V_{2}],$ and the corresponding contributions are still denoted by $\widetilde{F}(\cdots)$ for convenience.
@@
$\widehat{F}(w_{1},w_{2})$ $=\frac{1}{4\pi^{2}}\iint_{\mathbf{R}^{2}}\widetilde{F}(it_{1},it_{2})w_{1}^{-it_{1}}w_{2}^{-it_{2}}\mathrm{d}t_{1}\mathrm{d}t_{2},$
@@
$\widetilde{F}(s_{1},s_{2})$ $=\iint_{(\mathbf{R}^{+})^{2}}\widehat{F}(w_{1},w_{2})w_{1}^{s_{1}-1}w_{2}^{s_{2}-1}\mathrm{d}w_{1}\mathrm{d}w_{2}.$
@@
×$\ \ \ \ \times\iint_{\mathbf{R}^{2}}\widetilde{F}(it_{1},it_{2})\Big{(}\frac{n_{2}r}{m_{1}y_{1}}\Big{)}^{it_{1}}\Big{(}\frac{m_{1}r}{y_{2}n_{2}}\Big{)}^{it_{2}}\mathrm{d}t_{1}\mathrm{d}t_{2}.$
@@
×$\ \ \ \ \times\sum_{1\leqslant|h|\leqslant H}\sum_{|r|\sim R^{\prime}}\mathrm{e}(-hry)r^{i(t_{1}+t_{2})-1}$
×$\ \ \ \ \times\mathop{\sum\sum\sum\sum}_{\begin{subarray}{c}m_{1}\sim A_{1},~{}n_{1}\sim B_{1},~{}m_{2}\sim A_{2},~{}n_{2}\sim B_{2}\\ (m_{1}m_{2}n_{1}n_{2},q)=(m_{1}n_{1},m_{2}n_{2})=(m_{1},d_{2})=(m_{2},d_{2}^{\prime})=1\end{subarray}}\alpha^{\prime}_{d_{1}m_{1}}\alpha^{\prime\prime}_{d_{2}n_{1}}\beta^{\prime}_{d_{1}^{\prime}m_{2}}\beta^{\prime\prime}_{d_{2}^{\prime}n_{2}}$
×$\ \ \ \ \times\Big{(}\frac{n_{2}}{m_{1}}\Big{)}^{i(t_{1}-t_{2})}\mathrm{e}\Big{(}(m_{1}m_{2}y_{1}+n_{1}n_{2}y_{2})y\delta-\frac{qy_{2}}{m_{1}m_{2}}\Big{)}\mathrm{e}\Big{(}\frac{(q/\delta)hr\overline{m_{1}n_{1}}}{m_{2}n_{2}}\Big{)}.$
@@
$(a,h,m,n,r,s)\leftarrow(q/\delta,k,m_{1},n_{1},m_{2},n_{2}),\ \ \ (H,M,N,R,S)\leftarrow(R^{\prime}H,A_{1},B_{1},A_{2},B_{2}),$
@@
$\mathcal{R}^{*}_{2}$ $\ll(qNL)^{\varepsilon}\eta^{4}L(RHN)^{\frac{1}{2}}((RHN_{1}N_{2})^{\frac{1}{4}}+(N_{1}N_{2})^{\frac{1}{2}})$
×$\ \ \ \ \times\Bigg{[}q^{\vartheta}(N_{1}N_{2})^{\frac{1}{2}}\Big{(}1+\frac{qRH}{N^{2}}\Big{)}^{\frac{1}{2}}\Big{(}1+\frac{qRH}{N^{2}}+\frac{RH}{N_{1}N_{2}}\Big{)}^{\frac{1}{2}}\Bigg{]}^{\frac{1}{2}}+\eta^{4}NL(RH/N_{1}N_{2})^{\frac{1}{2}+\varepsilon}$
@@
Moreover, in the settings of Theorem 2.1, the coefficients $-β\boldsymbol{\alpha},\bbeta$ are arbitrary, and we appeal to Lemma 2.3 instead to bound the sums with Kloosterman fractions in (7.9). To this end, one may argue as above to group the variables, and the details are omitted here.
@@ 8.1. Twisted fourth moment of Dirichlet $L$-functions
As a (weaker) substitute of the Lindelöf Hypothesis, it is a reasonable approach to consider moments of Dirichlet $L$-functions. It has been known for a long time that
@@
for any $\varepsilon>0$ and $t\in\mathbf{R}$. In addition, we may also expect, for an arbitrary coefficient $\boldsymbol{\alpha}=(\alpha_{m})$ and any $\varepsilon>0$, that
@@
Recall that we have utilized (8.1) in the proof of Theorem 1.5. If one was allowed to use (8.2) with smooth and squarefree $q$, it would be possible to obtain the desirable constant $2$ in Theorem 1.5 in a larger range of $\varpi$.
@@ 8.2. Brun–Titchmarsh theorem on average
Hooley [Ho72] was the first who was able to study upper bounds of $\pi(x;q,a)$ non-tivially through extra averaging over $q$. This was largely motivated by the study of greatest prime factors of shifted primes, as approximations to the twin prime conjecture. For an individual $q$, one may also consider the average over the residue class $a\,(\operatorname{mod}{q})$. See also [Mo74, Ho75, Mo79, Iw82, Fo84b, Fo85, BH96, FR22] for these developments.
@@
In his ETH thesis, Löffel [Lo16] generalized Lemma 3.3, as well as other works of Fouvry, Kowalski and Michel [FKM14, FKM15], to squarefree moduli. This also allows us to relax $q$ in Theorem 1.2 to be squarefree. It is also possible to generalize Theorem 1.2 to arbitrary moduli along with the same line. As one may see from the proof of Lemma 3.3, it is necessary to evaluate averages of products of four Kloosterman sums to general moduli.
@@ 8.4. Twisted second moment of character sums
with $|c_{m}|\ll m^{\varepsilon}.$ We expect that $\mathcal{Z}\ll MLq^{\varepsilon}$ holds with $M$ as large as possible, building on the squareroot cancellation philosophy.
@@
Recall that Theorem 2.1 and Lemma 2.5 are proven with the aid of Lemma 2.3 on trilinear forms wiht Kloosterman fractions. However, Theorem 2.2 and Lemma 2.6 rely on the control of the quintilinear form in (7.1):
$-β\displaystyle\mathcal{Q}_{a}(\boldsymbol{\alpha},\bbeta)$ $=\sum_{h\leqslant H}\mathop{\sum_{m\leqslant M}\sum_{n\leqslant N}\sum_{r\leqslant R}\sum_{s\leqslant S}}_{(mn,rs)=(ns,a)=1}\alpha_{m,r}\beta_{h,n,s}\mathrm{e}\Big{(}\frac{ah\overline{mn}}{rs}\Big{)}.$
Due to the application of Kloostermania, the size of $a$ plays an important role in the settings of Theorem 2.2. Note that Lemma 2.3 follows from a careful application of amplification, so that only elementary means together with Weil’s bound for complete algebraic exponential sums are needed. A natural problem is to ask for an elementary bound for $-β\mathcal{Q}_{a}(\boldsymbol{\alpha},\bbeta)$ which is uniform in $a\neq 0$, and $-β\boldsymbol{\alpha},\bbeta$ are of convolution structures if necessary.
8.5. Hooley’s Hypothesis R and impacts
@@
are admissible in (1.3). Hypothesis R exhibits squareroot cancellations within incomplete Kloosterman sums. It seems far beyond the current technique to prove Conjecture 8.1 for general $q$, even in the case $|I|\approx q^{0.99}.$ On the other hand, Conjecture 8.1 corresponds to the exponent pair hypothesis in the theory of arithmetic exponent pairs, for which we hope $(\kappa,\lambda)=(0,\frac{1}{2})$ is admissible for (3.2). This illustrates that all of our previous arguments starting from additive characters should never beat (8.3) unconditionally even imposing factorization conditions on $q$.
However, as $\varpi\rightarrow 9/20-$ for instance, Theorem 1.5 shows that $C(\varpi)=5/(5-6\varpi)$ is admissible, as long as $q$ is squarefree and smooth. This is obviously sharper than the conjectural bound (8.3) which takes $C(\varpi)=6/(5-6\varpi).$ As one may see from previous sections, Theorem 1.5 is proved by using orthogonality of multiplicative characters, so that Lemma 3.6 on character sums has been applied to finish the proof. This phenomenon indicates that for $\varpi<1/2-$ it is clever to use multiplicative characters, and for $\varpi>1/2+$ additive ones might be more powerful instead. The exact transition range is very mysterious to the authors.

### v4

@@ 7.3. Estimates for exponential sums in $\mathcal{R}^{*}_{2}$
[…]
@@
We would like explore the cancellations among the exponentials $\mathrm{e}(\frac{(q/\delta)hr\overline{m_{1}n_{1}}}{m_{2}n_{2}})$ using Lemma 7.2. It is now necessary to separate variables. To do so, we introduce the Fourier transform of $F$:
$\widehat{F}(y_{1},y_{2})$ $=\iint_{\mathbf{R}^{2}}F(\lambda_{1},\lambda_{2})\mathrm{e}(-\lambda_{1}y_{1}-\lambda_{2}y_{2})\mathrm{d}\lambda_{1}\mathrm{d}\lambda_{2}.$
The Fourier inversion gives
$F(\lambda_{1},\lambda_{2})$ $=\iint_{\mathbf{R}^{2}}\widehat{F}(y_{1},y_{2})\mathrm{e}(\lambda_{1}y_{1}+\lambda_{2}y_{2})\mathrm{d}y_{1}\mathrm{d}y_{2}.$
Note that for all $\nu_{1},\nu_{2}\in\mathbf{N},$
[…]
@@
[…]
@@
which is also negligible. It thus suffices to treat the integral over $y_{1}\in I_{2}\cup I_{4}$. In a similar manner, we may also reduce the integral over $y_{2}$ to $V_{2},-U]\cup[U,V_{2}]$, where
@@
[…]
$\ \ \ \ \ \times\mathrm{e}\Big((m_{1}m_{2}y_{1}+n_{1}n_{2}y_{2})y\delta-\frac{qy_{2}}{m_{1}m_{2}}\Big)\mathrm{d}y_{1}\mathrm{d}y_{2}.$
@@
$(\varDelta\widehat{F})(w_{1},w_{2})$ $=\frac{1}{4\pi^{2}}\iint_{\mathbf{R}^{2}}\widetilde{F}(it_{1},it_{2})w_{1}^{-it_{1}}w_{2}^{-it_{2}}\mathrm{d}t_{1}\mathrm{d}t_{2},$
@@
$\widetilde{F}(s_{1},s_{2})$ $=\iint_{(\mathbf{R}^{+})^{2}}(\varDelta\widehat{F})(w_{1},w_{2})w_{1}^{s_{1}-1}w_{2}^{s_{2}-1}\mathrm{d}w_{1}\mathrm{d}w_{2}.$
@@
$\ \ \ \ \times\iint_{\mathbf{R}^{2}}\widetilde{F}(it_{1},it_{2})\Big(\frac{n_{2}r}{m_{1}y_{1}}\Big)^{it_{1}}\Big(\frac{m_{1}r}{y_{2}n_{2}}\Big)^{it_{2}}\mathrm{d}t_{1}\mathrm{d}t_{2}.$
@@
$\ \ \ \ \times\sum_{1\leqslant|h|\leqslant H}\sum_{|r|\sim R^{\prime}}\mathrm{e}(-hry)r^{i(t_{1}+t_{2})-1}$
$\ \ \ \ \times\mathop{\sum\sum\sum\sum}_{\begin{subarray}{c}m_{1}\sim A_{1},~n_{1}\sim B_{1},~m_{2}\sim A_{2},~n_{2}\sim B_{2}\\ (m_{1}m_{2}n_{1}n_{2},q)=(m_{1}n_{1},m_{2}n_{2})=(m_{1},d_{2})=(m_{2},d_{2}^{\prime})=1\end{subarray}}\alpha^{\prime}_{d_{1}m_{1}}\alpha^{\prime\prime}_{d_{2}n_{1}}\beta^{\prime}_{d_{1}^{\prime}m_{2}}\beta^{\prime\prime}_{d_{2}^{\prime}n_{2}}$
$\ \ \ \ \times\Big(\frac{n_{2}}{m_{1}}\Big)^{i(t_{1}-t_{2})}\mathrm{e}\Big((m_{1}m_{2}y_{1}+n_{1}n_{2}y_{2})y\delta-\frac{qy_{2}}{m_{1}m_{2}}\Big)\mathrm{e}\Big(\frac{(q/\delta)hr\overline{m_{1}n_{1}}}{m_{2}n_{2}}\Big).$
@@
$(a,h,m,n,r,s)$ $\leftarrow(q/\delta,k,m_{1},n_{1},m_{2},n_{2}),$
$(H,M,N,R,S,\xi)$ $\leftarrow(R^{\prime}H,A_{1},B_{1},A_{2},B_{2},(\log N)^{\frac{1}{3}}),$
@@
$\mathcal{R}^{*}_{2}$ $\ll(qNL)^{\varepsilon}\eta^{2}L\Big(\frac{\eta N^{3}}{q}\Big)^{\frac{1}{2}}\Big[\Big(\frac{\eta N^{2}N_{1}N_{2}}{q}\Big)^{\frac{1}{4}}+(N_{1}N_{2})^{\frac{1}{2}}\Big]\Big[\eta q^{\vartheta}(N_{1}N_{2})^{\frac{1}{2}}\Big(1+\frac{N^{2}}{qN_{1}N_{2}}\Big)^{\frac{1}{2}}\Big]^{\frac{1}{2}}$
$\ \ \ \ +\eta^{4}NL(\eta N^{2}/qN_{1}N_{2})^{\frac{1}{2}+\varepsilon}$
@@
[…]
@@ 8.1. Twisted fourth moment of Dirichlet $L$-functions
As a (weaker) substitute for the Lindelöf Hypothesis, it is a reasonable approach to consider moments of Dirichlet $L$-functions. It has been known for a long time that
@@
for any $\varepsilon>0$ and $t\in\mathbf{R}$. In addition, we may also expect, for any complex sequence $\boldsymbol{\alpha}=(\alpha_{m})$ and any $\varepsilon>0$, that
@@
Recall that we have utilized (8.1) in the proof of Theorem 1.5. If one were allowed to use (8.2) with smooth and squarefree $q$, it would be possible to obtain the desirable constant $2$ in Theorem 1.5 in a larger range of $\varpi$.
@@ 8.2. Brun–Titchmarsh theorem on average
Hooley [Ho72] was the first who was able to study upper bounds of $\pi(x;q,a)$ non-trivially through extra averaging over $q$. This was largely motivated by the study of greatest prime factors of shifted primes, as approximations to the twin prime conjecture. For an individual $q$, one may also consider the average over the residue class $a\,(\operatorname{mod}{q})$. See also [Mo74, Ho75, Mo79, Iw82, Fo84b, Fo85, BH96, FR22] for these developments.
@@
In his ETH thesis, Löffel [Lo16] generalized Lemma 3.3, as well as other works of Fouvry, Kowalski and Michel [FKM14, FKM15], to squarefree moduli. This also allows us to relax $q$ in Theorem 1.2 to be squarefree. It is also possible to generalize Theorem 1.2 to arbitrary moduli along the same lines. As one may see from the proof of Lemma 3.3, it is necessary to evaluate averages of products of four Kloosterman sums to general moduli.
@@ 8.4. Twisted second moment of character sums
with $|c_{m}|\ll m^{\varepsilon}.$ We expect that $\mathcal{Z}\ll MLq^{\varepsilon}$ holds with $M$ as large as possible, building on the square-root cancellation philosophy.
@@
Recall that Theorem 2.1 and Lemma 2.5 are proven with the aid of Lemma 2.3 on trilinear forms with Kloosterman fractions. However, Theorem 2.2 and Lemma 2.6 rely on the control of the quintilinear form in (7.1):
[…]
8.5. Hooley’s Hypothesis R and its impacts
@@
are admissible in (1.3). Hypothesis R exhibits square-root cancellations within incomplete Kloosterman sums. It seems far beyond the current techniques to prove Conjecture 8.1 for general $q$, even in the case $|I|\approx q^{0.99}.$ On the other hand, Conjecture 8.1 corresponds to the exponent pair hypothesis in the theory of arithmetic exponent pairs, for which we hope $(\kappa,\lambda)=(0,\frac{1}{2})$ is admissible for (3.2). This illustrates that all of our previous arguments starting from additive characters should never beat (8.3) unconditionally, even when imposing factorization conditions on $q$.
However, as $\varpi\rightarrow 9/20$- for instance, Theorem 1.5 shows that $C(\varpi)=5/(5-6\varpi)$ is admissible, as long as $q$ is squarefree and smooth. This is obviously sharper than the conjectural bound (8.3) which takes $C(\varpi)=6/(5-6\varpi).$ As one may see from previous sections, Theorem 1.5 is proved using the orthogonality of multiplicative characters, and hence Lemma 3.6 on character sums can be applied to complete the argument. This phenomenon indicates that for $\varpi<1/2$- it is clever to use multiplicative characters, and for $\varpi>1/2+$ additive ones might be more powerful instead. The exact transition range is very mysterious to the authors.